Sarkis is out looking for flour and has six markets on his list. They have different suppliers, so the presence of flour at one market is independent of all the others. The probability that any individual market has flour is 0.34, so what is the probability that at least one of the markets has flour?
Answer #1

The probability that any individual market has flour = 0.34

The probability that any individual market dies not have flour = 1 - 0.34 = 0.66

The probability that at least one of the six markets has flour = 1- The probability that none of the six markets has flour = 1 - = 1 - 0.0827 = 0.9173